OpenStudy (anonymous):

what is the range of 2*sin^-1(x)

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

range of arcsine is not minus infinity to infinity, it is \([-\frac{\pi}{2},\frac{\pi}{2}]\)

OpenStudy (anonymous):

@Sheng i dont think so. if inverse circular function is restricted

OpenStudy (anonymous):

isnt that domain?

OpenStudy (anonymous):

multiply by two to get your answer

OpenStudy (anonymous):

no that is the range

OpenStudy (anonymous):

@satellite73 i dont care about the 2 infront of arcsine?

OpenStudy (anonymous):

domain is \([-1,1]\) because that is the range of sine @ksw19372 yes you do care about the two

OpenStudy (anonymous):

that is why i said multiply by 2

OpenStudy (anonymous):

then how do i get [-pi/2,pi/2]?? i dont know why but i am keep getting [-pi/4,pi/4]

OpenStudy (anonymous):

if arcsine is \(-\frac{\pi}{2}\) then 2arcsine is \(-\pi\)

OpenStudy (anonymous):

i thought range was bound of y values?

OpenStudy (anonymous):

i think you are confusing domain and range

OpenStudy (anonymous):

@cheng its inverse

OpenStudy (anonymous):

domain of arcsine is the range of sine

OpenStudy (anonymous):

ooh! thx for pointing that out, inverse functions domain/range are flipped :)

OpenStudy (anonymous):

range of sine is \([-1,1]\) so domain of arcsine is \([-1,1]\)

OpenStudy (anonymous):

@satellite73 what about the range of sin^-1(x)?

OpenStudy (anonymous):

now to make sine one to one you restrict the domain to \([-\frac{\pi}{2},\frac{\pi}{2}]\) so that now becomes the range of arsine

OpenStudy (anonymous):

so range of arcsine is \([-\frac{\pi}{2},\frac{\pi}{2}]\) and range of \(2\arcsin(x)\) is \[[-\pi,\pi]\]

OpenStudy (anonymous):

since you multiply your answer by two

OpenStudy (anonymous):

so the answer of my question is [-pi,pi]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i get it !

OpenStudy (anonymous):

\[\arcsin(-1)=-\frac{\pi}{2}\] and so \[2\arcsin(-1)=-\pi\]

OpenStudy (anonymous):

ok good!

OpenStudy (anonymous):

@satellite73 thank you so much. please help in the further questions later on! :)